OpenStudy (anonymous):

The slope of the line whose equation is y - 3 = 0 is 0 3 no slope

OpenStudy (mertsj):

put the equation in y = mx + b form

OpenStudy (anonymous):

is it no slope

OpenStudy (mertsj):

Put the equation in y = mx+b form

OpenStudy (anonymous):

ax + by + c =0 m = -a/b so y - 3 = 0 0x + 1y - 3 =0 m = -0/1 = 0